# von Neumann entropy and quantum version of thermodynamic entropy

### v2

@@ IV.3 Subsystem of a quantum system
For our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
@@
To numerically calculate the VN entropy of the subsystem, we first find $\rho_{sb}$ from the state ($\rho$) of the full system: $\rho_{sb}=\text{Tr}_{B}(\rho)$. We then use the formula Eq. 3 to calculate the VN entropy of the subsystem.
To numerically calculate the corresponding TH entropy, consider that the full system is prepared with an energy lying in $(E,E+\Delta E]$. As discussed earlier, the system will be in a mixed state $\rho_{mc}$, as appear in Eq. 8. The TH (or qBN) entropy of the full system is $S_{TH}\equiv S_{qBN}=k_{B}\ln D$, as in Eq. 10, where $D$ is the number of eigenvalues of the system lying in the aforementioned energy shell. Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $k_{B}\ln D$ or to the $k_{B}\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$. Accordingly,
$S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto k_{B}\ln(DOS),$ (14)
@@
Now, for different cases, we calculate the VN entropy and the TH entropy of a subsystem to study their equivalence. First, we consider the case when the full system is in an eigenket |$\ket{n}$ of the Hamiltonian of Eq. 13. In this case the subsystem will be represented by the reduced density matrix $\rho_{sb}\equiv\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$. Now using Eq. 3, we can calculate the VN entropy of the subsystem when the full system is in the eigenket |$\ket{n}$:
$S^{sb}_{VN}=-k_{B}\text{Tr}(\rho_{sb}\ln\rho_{sb}).$ (15)
@@
The values of $S^{sb}_{VN}$ for different eigenkets across the spectrum can be found in Fig. 2 (for $\Delta_{2}=0.0,0.2$ and $0.5$). We also show the corresponding logarithm of the density-of-states (DOS) for the full system. As mentioned earlier, when the system does not thermalize ($\Delta_{2}=0.0$), the individual reduced density matrices are quite different from each other (even within a narrow energy shell) and accordingly, $S^{sb}_{VN}$ takes quite different values as one goes from one eigenket to another within the energy shell. In this case the VN entropy of the subsystem is not, in general, proportional to its qBN entropy, i.e., to $\ln(DOS)$. On the other hand, when the system thermalizes (or the ETH holds), $S^{sb}_{VN}$ takes very close values as one goes from one eigenket to another within an energy shell (see, for example, results for $\Delta_{2}=0.5$). In Fig. 3, we plot the average VN entropy ($\overline{S}^{sb}_{VN}$) of the subsystem against the logarithm of DOS (individually for left and right halves of the spectrum). We take the average of $S^{sb}_{VN}$ over a small energy shell to get better numerical results. The linearity of the plots shows that the VN entropy of a subsystem is equivalent to its TH (or qBN) entropy, i.e., $S^{sb}_{VN}\approx\overline{S}^{sb}_{VN}\equiv S^{sb}_{qBN}$.
Next, we consider the case when the full system is in a mixed state ($\rho_{mc}$), as appears Eq. 8 (with $p_{n}=\frac{1}{D}$). In this case, the subsystem will be in the following mixed state,
$\rho_{sb}=\frac{1}{D}\sum_{n}\rho_{sb}^{(n)},$ (16)
where the reduced density matrix $\rho_{sb}^{(n)}=\text{Tr}_{B}(\ket{n}\bra{n})$ is calculated for a single eigenket |$\ket{n}$ of the full system. Here the sum is over the all eigenkets from a small (fixed) energy shell. There are $D$ eigenkets in the shell. Using Eq. 15, we now can calculate the VN entropy of the subsystem directly from $\rho_{sb}$ of Eq. 16. The value of $S^{sb}_{VN}$ across the energy spectrum can be found in Fig. 4.
Figure 4: The subsystem VN entropy ($S^{sb}_{VN}$) when the full system is in the microcanonical density matrix $\rho_{mc}$. Here, the system (subsystem) size is 16 (6).
The plot of $S^{sb}_{VN}$ against $\ln(DOS)$ shows linearity, similar to what we find in Fig. 3. We note that, as expected, we get the said linear plot even when the system is in integrable regime ($\Delta_{2}=0.0$).
@@
It may be worth mentioning here that, the result that the subsystem’s average VN entropy is proportional to system’s $\ln(DOS)$, is earlier reported in [36] in a different context. We also note that, elsewhere [37], the TH entropy of a subsystem is defined as what we call here the VN entropy of a subsystem ($S^{sb}_{VN}$) when the full system is in $\rho_{mc}$. As discussed in this work, the TH entropy of a subsystem is appropriately calculated as the logarithm of $DOS$ (since, $S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS)$ under some general conditions).
V Conclusion
@@
In this work we analyze whether and when the von Neumann (VN) entropy is equal or equivalent to the thermodynamic (TH) entropy. In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly. We find that, when a large isolated system is in a mixed state (described by microcanonical ensemble), its VN entropy is equivalent to its TH entropy. When the system-bath combination is in a mixed state, or in a typical pure state, the VN entropy of the (open) system is equivalent to its TH entropy. This result also holds for a subsystem (of a much larger system) but the analytical demonstration is not straightforward. Through numerical analysis, we show that the VN entropy of a subsystem is equivalent to its TH (qBN) entropy when the full system (subsystem-bath combination) is described by a microcanonical ensemble or when the full system is in a typical pure state. In this work we also numerically demonstrate that, for the states that are atypical (or for which the ETH does not hold), the VN entropy and the TH entropy of a subsystem are not the same.
Acknowledgements.

### v3

@@ 4.4 Numerical results
For all our numerical results, we take the following one-dimensional (with open boundary condition) Heisenberg spin-1/2 chain with the next-nearest neighbor interactions:
@@
To numerically calculate the TH entropy, consider that the full system is prepared with an energy lying in $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$. The TH (or qBN) entropy of the full system is $S_{TH}\equiv S_{qBN}=\ln d_{E}$, as in Eq. 10, where $d_{E}$ is the number of eigenvalues of the system lying in the aforementioned energy shell. Now since the TH entropy is additive, we can assume that the TH entropy of the subsystem (denoted as $S^{sb}_{TH}$ ) is proportion to $\ln d_{E}$ or to the $\ln(DOS)$, where $DOS$ is the density-of-states of the full system at the energy scale $E$. Accordingly,
$S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS),$ (21)
@@
Now while numerically calculating the VN entropy of the subsystem, we consider two separate cases - full system in a mixed state and full system in a pure state. First consider that the full system is in mixed state $\rho_{mc}$, as appear in Eq. 8. Let $\rho_{sb}$ be the corresponding reduced density matrix of the subsystem: $\rho_{sb}=\mathrm{Tr}_{B}(\rho_{mc})$. More specifically, $\rho_{sb}=\frac{1}{d_{E}}\sum^{\prime}_{n}\rho_{sb}^{(n)}$, where the reduced density matrix $\rho_{sb}^{(n)}=\mathrm{Tr}_{B}(\ket{n}\bra{n})$ is calculated for a single eigenket |$\ket{n}$ of the full system. Here the primed sum is over the all eigenkets associated with a small (fixed) energy shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$. There are $d_{E}$ eigenkets associated with the shell. The VN entropy of the subsystem, then, would be $S^{sb}_{VN}=-\mathrm{Tr}_{sb}(\rho_{sb}\ln\rho_{sb})$. Thus, in the current scenario, we have the following:
$\rho_{sb}$ $=$ $\frac{1}{d_{E}}{\sum_{n}}^{\prime}\rho_{sb}^{(n)},~{}\mathrm{with}~{}\rho_{sb}^{(n)}=\mathrm{Tr}_{B}(\ket{n}\bra{n})$
$S^{sb}_{VN}$ $=$ $-\mathrm{Tr}_{sb}(\rho_{sb}\ln\rho_{sb}).$ (22)
The plot of $S^{sb}_{VN}$ across the energy spectrum of the full system can be found in Fig. 4. The plot of $S^{sb}_{VN}$ vs. $\ln(DOS)$ at different energy scale ($E$) is linear and very similar to the plot of $\overline{S}^{sb}_{VN}$ vs. $\ln(DOS)$ for the thermalized system as discussed below (and the corresponding plot can be found in Fig. 3). The linearity confirms that the VN entropy and the TH entropy (or qBN entropy) of a subsystem are equivalent when the full system is in a mixed state (described by a microcanonical ensemble).
Next, we consider the case where the full system is in a pure state. Once again, when the full system is described by a pure-state density matrix $\rho$, the reduced density matrix of the subsystem would be: $\rho_{sb}=\mathrm{Tr}_{B}(\rho)$. Consequently, the VN entropy of the subsystem is $S^{sb}_{VN}=-\mathrm{Tr}_{sb}(\rho_{sb}\ln\rho_{sb})$. Here we calculate this VN entropy corresponding to each eigenket from the pre-specified energy shell, and then find the average VN entropy $\overline{S}^{sb}_{VN}$. In the current scenario, we have the following:
$\rho^{(n)}_{sb}$ $=$ $\mathrm{Tr}_{B}(\ket{n}\bra{n})$
$\overline{S}^{sb}_{VN}$ $=$ $-\frac{1}{d_{E}}{\sum_{n}}^{\prime}\mathrm{Tr}_{sb}(\rho^{(n)}_{sb}\ln{\rho^{(n)}_{sb}}),$ (23)
where the primed sum is over all the eigenkets associated with the shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$.
@@
The values of $S^{sb}_{VN}$ for the individual eigenkets across the spectrum can be found in Fig. 2 (for $\Delta_{2}=0.0,0.2$ and $0.5$). We also show the corresponding logarithm of the density-of-states (DOS) for the full system. As mentioned earlier, when the system does not thermalize ($\Delta_{2}=0.0$), the eigen spectra of individual reduced density matrices are quite different from each other (even within a narrow energy shell) and accordingly, $S^{sb}_{VN}$ takes quite different values as one goes from one eigenket to another within the energy shell. In this case the VN entropy of the subsystem is not, in general, proportional to its qBN entropy, i.e., to $\ln(DOS)$ with proper scaling factor.
On the other hand, when the system thermalizes (or the ETH holds), $S^{sb}_{VN}$ takes very close values as one goes from one eigenket to another within an energy shell (see, for example, results for $\Delta_{2}=0.5$). In Fig. 3, we plot the average VN entropy ($\overline{S}^{sb}_{VN}$ as in Eq. 4.4) of the subsystem against the logarithm of DOS (separately for the left and right halves of the spectrum). The linearity of this plot is as per our prediction in the last subsection (Sec. 4.3), and it confirms that the VN entropy and the TH entropy (or qBN entropy) of a subsystem are equivalent when the full system is in pure state (this holds when the full system thermalizes).
We conclude this subsection with a few remarks. When a system is thermalized, the VN entropy of its subsystem, corresponding to any individual eigenket from a narrow energy shell, is quite close to its mean value. This is evident in Fig.2(a). For any thermalized or non-thermalized system, when we average the reduced density matrices associated with a narrow energy shell, the eigen spectrum of this resultant density matrix is expected to be very similar to the eigen spectrum of a reduced density matrix associated with a thermalized system. As a consequence, the results from Eq. 4.4, irrespective of the type of system, will be very similar to the results from Eq. 4.4 when the system is thermalized. This is evident if one compares Fig. 2(a) and Fig. 4.
@@
We also remark here that, in a different context [28], numerically it was previously observed that the subsystem’s average VN entropy is proportional to the system’s $\ln(DOS)$. We also note that, elsewhere [6], the TH entropy of a subsystem is defined as what we call here the VN entropy of a subsystem ($S^{sb}_{VN}$) when the full system is in a mixed state ($\rho_{mc}$). As discussed in this work, the TH entropy of a subsystem is appropriately calculated as the logarithm of $DOS$ (since, $S^{sb}_{TH}\equiv S^{sb}_{qBN}\propto\ln(DOS)$ under some general conditions).
5 Conclusion
@@
In this work we analyze whether and when the von Neumann (VN) entropy is equal or equivalent to the thermodynamic (TH) entropy. In reviewing the major objections about the VN entropy, we explain that the concerns about the time-invariance and subadditivity properties can be addressed convincingly. We separately address the question of equivalence for isolated system, open system, and subsystem (both when the full system is in a mixed state or in a pure state). In this context, we show that the result of Page and Sen can be strengthened for a thermalized system, and further argued that this new result implies the equivalence of the VN entropy and the TH entropy of a subsystem when the full system is in a (random) pure state associated with a narrow energy shell. We provide numerical data to demonstrate or assert some of our key results.
6 Acknowledgement